A loan is made for five years at a simple interest rate of 12% per annum. What is the equivalent annual effective rate of discount during the fourth year of the loan?
Answer #1

Here i = 0.12 and n = 4

Thus effective rate of discount = i/(1+i(n-1))

= 0.12/(1+0.12*(4-1))

= 0.12/1.36

= 8.82%